heat produced in a conductor of resistance 4.2ohm with 10A current is 

Grade:8

3 Answers

Umakant biswal
7 years ago
@ gayathri 
I think the time must be given in the question , if not then u can directly apply the joules law of heating in that case 
which formula is 
H = I SQUARE R 
SO, BY CALCULATING 
H = 10 square * 4.2 
H= 100 * 4.2 
H= 420 JOULE . 
hope it clears your doubt 
all the best ...
haripriya
7 years ago
heat produses many ways like sunheat is very product it can used elevators solar many types arein  our sun heat
 
PRABHU KIRAN
3 years ago
We know that,H=I2RT
As we know it from JOULES LAW,
But T is not given,so,H=I2R
H=(10*10*4.2)
H=420JOULES/SEC
H=420/(4.2)=100CAL/SEC